# Joint User Association and Resource Allocation for Multi-Cell Networks with Adaptive Semantic Communication

### 05 Jan 2024 v2

@@ III-D Problem Formulation
In this paper, we aim to maximize the total utility by jointly optimizing the user association and resource allocation under energy consumption, wireless bandwidth, and completion time constraints. Based on the models described above, the corresponding optimization problem can be formulated as follows:
@@
where $\bm{x},\bm{z},\bm{f},\bm{P},\bm{c},\bm{d}$ are vetors of corresponding decision variables. Constraint (10a) contains the user association and RB allocation constraints described above. Constraint (10b) ensures the semantic information transmission is completed before the maximum delay $T^{max}$. Constraint (10c) requires the total energy consumption of WD $n$ does not exceed a given threshold $E^{max}_{n}$. Constraints (10d), (10e), (10f), and (10g) define the ranges of feasible values for the decisions variables, where $f^{max}_{n}$ and $P^{max}_{n}$ are the maximum CPU frequency and transmission power of WD $n$, respectively.
The formulated problem is Mixed-Integer Programming (MIP) and can be hard to solve. Specifically, we will show that our problem is strongly NP-hard, so it does not even have a pseudo-polynomial time algorithm. Before stating our result, we give the definition of binary utility functions: we say an utility function $u^{a}(c,d)$ is binary if $u^{a}(c,d)=u^{a}_{0}$ for all $(c,d)\geq(c^{a}_{0},d^{a}_{0})$ and $u^{a}(c,d)=0$ otherwise, where $u^{a}_{0},c^{a}_{0},d^{a}_{0}$ are positive constants.
@@
Problem (10) is strongly NP-hard even if (i) there is only one WD for each running application and its utility function is binary, (ii) the energy budget $E^{max}_{n}$ is sufficiently large, and (iii) the channel gain $h^{m}_{n}$ and interference $I_{m}$ for different BS $m$ is equal.
Proof.
Please see Appendix A. ∎
Thereom 1 indicates that problem (10) is hard to solve even for a significantly simplified case. Therefore, in general, it is computationally intractable to obtain the exact solution of problem (10). In the next section, we will design an approximate algorithm that produces near-optimal solutions for problem (10) in polynomial time.
@@ IV Algorithm Design and Performance Analysis
In this section, we will progressively solve the formulated problem through three stages. Firstly, we find the optimal scheduling decisions $\bm{f},\bm{P},\bm{c},\bm{d}$ when the user association and RB allocation are given. Based on this result, we can solve the optimal RB allocation under any user association. Finally, we can obtain the optimal user association by employing a relaxation-then-refinement approach.
@@ IV-A Scheduling Subproblem for Each WD
If the user association $\bm{x}$ and RB allocation $\bm{z}$ are given, problem (10) can be decomposed into $N$ subproblems concerning the scheduling of each WD. Specifically, for any WD $n\in\mathcal{N}_{a}$ associated with BS $m$, its scheduling subproblem can be formulated as:
@@
We say a computation and communication workload pair $(c_{n},d_{n})$ is achievable if there exists $f_{n},P_{n}$ such that constraints (11a)-(11c) are satisfied. The set of all achievable workload pairs is referred to as the achievable region. Since $u^{a}(c_{n},d_{n})$ is non-decreasing, the optimal workload pair $(c^{*}_{n},d^{*}_{n})$ must lie on the boundary of the achievable region, i.e. there is no achievable $(c_{n},d_{n})$ such that $c_{n}=c^{*}_{n}$ and $d_{n}>d^{*}_{n}$. To obtain the achievable region, we only need to calculate the maximum achievable $d_{n}$ for any given $c_{n}$, denoted by $d^{max}_{n}(c_{n})$. The calculation of $d^{max}_{n}(c_{n})$ can be formulated as another optimization problem
@@
To solve problem (12), we first show that its optimal solution satisfies the following condition.
@@
Proof.
Please see Appendix B. ∎
According to Theorem 2, we can replace constraint (11a) and (11b) with (13) and (14). Substituting (2) and (6) into (13) and recalling that WD $n$ is associated with BS $m$ (i.e. $x^{m}_{n}=1$) result in
@@
Combining (15) with (11c) we get the feasible region of $f_{n}$:
@@
Hence, problem (12) is equivalent to finding the optimal $f_{n}$ that maximizes (17) subject to (16), which can be solved by letting $\frac{\partial d_{n}}{\partial f_{n}}=0$ and comparing the values of $d_{n}$ achieved by every critical point in the feasible region. Clearly, the optimal $f_{n}$ is a function of $c_{n}$, denoted by $f^{*}_{n}(c_{n})$. According to (17), we can regard $d_{n}$ as a function of $c_{n}$ and $f_{n}$. Then we have $d^{max}_{n}(c_{n})=\max_{f_{n}}d_{n}(c_{n},f_{n})=d_{n}(c_{n},f^{*}_{n}(c_{n}))$. Unfortunately, the expression of $\frac{\partial d_{n}}{\partial f_{n}}=0$ is fairly complex so we are unable to derive an analytical expression for $f^{*}_{n}(c_{n})$. However, we can still calculate the value of $f^{*}_{n}(c_{n})$ through numerical methods.
Once we know how to calculate $d^{max}_{n}(c_{n})$, problem (11) is equivalent to $\max_{c_{n}}u^{a}(c_{n},d^{max}_{n}(c_{n}))$. Similarly, the optimal $c_{n}$ can be obtained by letting
@@
$\quad=\frac{\partial u^{a}}{\partial c_{n}}+\frac{\partial u^{a}}{\partial d^{max}_{n}}\times\frac{\partial d_{n}}{\partial c_{n}}=0$
where the second equality holds because $d^{max}_{n}(c_{n})=d_{n}(c_{n},f^{*}_{n}(c_{n}))$ and the third equality holds because $\frac{\partial d_{n}}{\partial f^{*}_{n}}=0$. Similarly, the optimal $c_{n}$ can only be solved by numerical methods. After obtaining the optimal $c_{n}$, the optimal values of the rest decision variables can be calculated easily.
IV-B RB Allocation Subproblem for Each BS
In the previous subsection, we have obtained the optimal $\bm{f},\bm{P},\bm{c},\bm{d}$ for any given $\bm{x}$ and $\bm{z}$. In this subsection, we still assume $\bm{x}$ is given and show how to find the optimal $\bm{z}$. Let $u^{*}_{n}(z^{m}_{n})$ be the optimal objective value of problem (11). The meaning of $u^{*}_{n}(z^{m}_{n})$ is the maximum achievable utility of WD $n$ when $z^{m}_{n}$ RBs are allocated to it. Since $\bm{x}$ is given, we can use $\mathcal{N}_{m}$ to denote the set of WDs associated with BS $m$. Then the RB allocation subproblem for BS $m$ becomes
@@ IV-B RB Allocation Subproblem for Each BS
In this subsection, we will design two different algorithms for problem (18) based on the properties of the utility function $u^{a}(c,d)$. In practice, the most commonly used utility function is the result accuracy, e.g. the classification accuracy in image recognition tasks. In this case, the utility function usually exhibits concavity and other properties that can be exploited to accelerate the solving of problem (18). However, in some scenarios, we may encounter more general utility functions. Therefore, we also design an algorithm that does not require the concavity of the utility function.
@@ IV-B1 RB Allocation with Concave Utility
If the utility function is the result accuracy or a concave function of the result accuracy, then it should satisfy the following three properties. Firstly, according to the results in [22, 23, 24, 25, 18, 19, 20, 21], the marginal return of $c$ and $d$ is diminishing. Hence, we can assume:
@@
Secondly, the marginal return of computation workload should be non-decreasing with respect to the amount of transmitted data. This is because the improvement in result accuracy due to increased computation needs to be propagated to the destination through data transmission. If the amount of transmitted data is limited, then the enhancement obtained at the destination will be discounted. An extreme example is the marginal return of $c$ is 00 when $d=0$, as $u^{a}(c,0)=0$ for all $c$. Therefore, we have:
@@
Assumption 3 requires $\frac{\partial u^{a}}{\partial d}$ decreases at least at a linear rate of $d$. A common function that satisfies this assumption is the logarithmic function. In practice, this assumption generally does not hold strictly, but it can be regarded as approximately valid within the range of values that we are interested in.
With the above assumptions, we can proceed to demonstrate the concavity of $u^{*}_{n}(z^{m}_{n})$. However, $z^{m}_{n}$ is a discrete variable and the concavity for discrete functions is not well-defined. Therefore, we relax $z^{m}_{n}$ to a real-value variable $\hat{z}^{m}_{n}$ and prove that $u^{*}_{n}(\hat{z}^{m}_{n})$ is concave.
@@
Proof.
Please see Appendix C. ∎
Due to the concavity of $u^{*}_{n}(\hat{z}^{m}_{n})$, to solve (18), we only need to greedily allocate RBs to the WD that contributes most to the increase of utility. The detailed steps are as follows. At the beginning, we initialize $z^{m}_{n}=0$ for all $n\in\mathcal{N}_{m}$. After that, for each WD $n\in\mathcal{N}_{m}$, we calculate the increase of utility if an extra RB is allocated to it, i.e. $\Delta u_{n}=u^{*}_{n}(z^{m}_{n}+1)-u^{*}_{n}(z^{m}_{n})$. If WD $n^{\prime}$ has the greatest increase of utility, i.e. $n^{\prime}=\operatorname*{arg\,max}_{n\in\mathcal{N}_{m}}\Delta u_{n}$, we allocate one extra RB to it and update $z^{m}_{n^{\prime}}=z^{m}_{n^{\prime}}+1$. This process continues until all of the $K_{m}$ RBs are allocated.
Our algorithm is summarized in Algorithm 1. Notice that we only modify the values of $z^{m}_{n}$ for one WD at each iteration. Therefore, for the loop in line $3$-$5$, we only need to calculate $\Delta u_{n}$ for all $n\in\mathcal{N}_{m}$ for the first time we entered the loop. After that, we only need to update $\Delta u_{n}$ for the WD whose $z^{m}_{n}$ is modified in the previous iteration. Apparently, the time complexity of the whole algorithm is $O(|\mathcal{N}_{m}|K_{m})$, where $|\mathcal{N}_{m}|$ is the number of elements in $\mathcal{N}_{m}$.
@@ IV-B2 RB Allocation with General Utility
For general utility functions, we cannot leverage the concavity of $u^{*}_{n}(\hat{z}^{m}_{n})$ to facilitate algorithm design. However, we can still solve the RB allocation subproblem (18) through dynamic programming. Specifically, let $u_{m}(k,j)$ be the maximum total utility of BS $m$ if there are only $k$ available RBs and $j$ WDs in BS $m$, i.e. $u_{m}(k,j)$ is the optimal objective value of the following optimization problem:
@@
Apparently, the optimal objective value of problem (18) equals $u_{m}(K_{m},|\mathcal{N}_{m}|)$. For convenience, we use $z_{m}(k,j)$ to denote the optimal RB allocation for the last WD (i.e. WD $j$) in (19). Clearly, when there is only one WD in the system, we should allocate all RBs to it. Therefore, we have
$u_{m}(k,1)=u^{*}_{1}(k)\mbox{ and }z_{m}(k,1)=k.$
Otherwise, for any $j\in\{2,3,\dots,N\}$, the value of $u_{m}(k,j)$ can be calculated through
$u_{m}(k,j)=\max_{z^{m}_{j}}\{u_{m}(k-z^{m}_{j},j-1)+u^{*}_{j}(z^{m}_{j})\}.$ (20)
Correspondingly, $z_{m}(k,j)$ is the optimal $z^{m}_{j}$ that maximizes (20). Based on (20), we can finally obtain $u_{m}(K_{m},|\mathcal{N}_{m}|)$. Then the optimal $z^{m}_{n}$ can be obtained by backtracking the optimal allocation decisions recorded in $z_{m}(k,j)$. Specifically, we can let $z^{m}_{|\mathcal{N}_{m}|}=z_{m}(K_{m},|\mathcal{N}_{m}|)$ and $z^{m}_{j}=z_{m}(K_{m}-\sum_{n=j+1}^{|\mathcal{N}_{m}|}z^{m}_{n},j)$ for $j\in\{1,\dots,|\mathcal{N}_{m}|-1\}$.
The detailed steps are summarized in Algorithm 2. When calculating $u_{m}(k,j)$ using (20), we need to iterate over all possible values of $z^{m}_{j}$, which amount to a total of $k$ values. Then one can easily verify that the time complexity of Algorithm 2 is $O\left(|\mathcal{N}_{m}|K^{2}_{m}\right)$.
@@
1:  Initialization: $u_{m}(k,1)=u^{*}_{1}(k),\ z_{m}(k,1)=k,\ \forall k\in\{1,\dots,K_{m}\}$;
@@
4:        Calculate $u_{m}(k,j)$ according to (20) and record the corresponding $z_{m}(k,j)$.

### 09 Mar 2026 v3

@@ III-D Problem Formulation
In this paper, we aim to maximize the overall system utility by jointly optimizing user association, resource allocation, and SemCom scheme, subject to constraints on energy consumption, spectrum availability, and delay. Based on the system models developed above, the corresponding optimization problem can be formulated as follows:
@@
Here, $\bm{x},\bm{z},\bm{f},\bm{P},\bm{c},\bm{d}$ are the vectors of decision variables. Constraint (8a) contains the user association and RB allocation constraints described above. Constraint (8b) ensures that the total latency, comprising both computation and transmission time, does not exceed the maximum delay threshold $T^{max}$. Constraint (8c) limits the total energy consumption of each WD to a specified maximum $E^{max}_{n}$. Constraints (8d) and (8e) define the feasible ranges for CPU frequency and transmission power, bounded by their respective hardware limits $f^{max}_{n}$ and $P^{max}_{n}$. Finally, constraints (8f) and (8g) define the domain of binary user association variables and integer-valued RB allocation variables.
The formulated optimization problem is a mixed-integer programming (MIP) problem, which is generally challenging to solve. In the literature, NP-hardness is commonly employed to characterize the inherent difficulty of such problems. However, establishing NP-hardness alone does not necessarily imply that a problem is intractable in practice. In particular, weakly NP-hard problems—such as the classical 0-1 knapsack problem—can be addressed using pseudo-polynomial time algorithms, and are often solved efficiently through dynamic programming [22]. In contrast, we demonstrate that the problem under consideration is strongly NP-hard, which indicates it admits no pseudo-polynomial time algorithm and is fundamentally more complex to solve.
@@
Problem (8) is strongly NP-hard even if (i) there is only one WD for each running application, (ii) the energy budgets $E^{max}_{n}$ are sufficiently large, and (iii) the channel gains $h^{m}_{n}$ and interference levels $I_{m}$ are identical across all SBSs.
Proof:
We consider a specific class of two-valued utility functions, defined as $u^{a}(c,d)=u^{a}_{0}$ for all $(c,d)\geq(c^{a}_{0},d^{a}_{0})$ and $u^{a}(c,d)=0$ otherwise, where $u^{a}_{0},c^{a}_{0},d^{a}_{0}$ are application-specific positive constants. According to condition (i), when WD $n$ is specified, the corresponding application $a$ is also determined. For notational convenience, we omit the superscript $a$ and use $u_{n}$ to denote $u^{a}_{0}$. Since the utility function is two-valued, WD $n$’s utility is either 0 or $u_{n}$. If the utility is 0, there is no need to associate it with any SBS. Therefore, without loss of generality, we can assume $c_{n}=c^{a}_{0}$, $d_{n}=d^{a}_{0}$, and WD $n$ takes utility $u_{n}$ if it is associated with some SBS $m$. Then the objective function of problem (8) can be re-written as $\sum_{n\in\mathcal{N}}\sum_{m\in\mathcal{M}}x^{m}_{n}u_{n}$.
According to condition (ii), the energy constraint (8c) is redundant, allowing us to fix the CPU frequency and transmission power at their maximum values, i.e., $f_{n}=f^{max}_{n}$ and $P_{n}=P^{max}_{n}$. Since both $f_{n}$ and $c_{n}$ are fixed, the computation time $T^{c}_{n}$ becomes a constant. Therefore, the delay constraint (8b) is equivalent to $T^{t}_{n}\leq T^{max}-T^{c}_{n}$, where the right-hand side is a constant. By substituting the expression for $T^{t}_{n}$ from (6), this constraint can be interpreted as a lower bound on the number of RBs allocated to WD $n$. Consequently, combining with condition (iii), the delay constraint (8b) can be equivalently expressed as: $z^{m}_{n}\geq x^{m}_{n}C_{n}$, where $C_{n}$ is a constant independent of $m$. Combining with constraints (5) and (8g), we derive an equivalent constraint on $x^{m}_{n}$:
where $\lceil\cdot\rceil$ is the ceiling function that returns the smallest integer which is greater than or equal to the input value. As a result, the original problem (8) reduces to a classical 0-11 multiple knapsack problem, which is known to be strongly NP-hard [22]. ∎
Theorem 1 highlights the intrinsic complexity of problem (8), showing that it remains computationally intractable even in highly simplified settings. Consequently, it is generally infeasible to obtain the exact optimal solution for non-trivial network scale. To address this issue, we propose an efficient algorithm capable of producing near-optimal solutions in polynomial time.
@@ IV Algorithm Design and Performance Analysis
In this section, we solve the formulated optimization problem through a three-stage framework. First, we determine the optimal scheduling variables $\bm{f},\bm{P},\bm{c},\bm{d}$ assuming fixed user association and RB allocation. Building on this result, we can optimize the RB allocation for any given user association. Finally, we derive the optimal user association by applying a relaxation-and-refinement strategy.
@@ IV-A Scheduling Subproblem for Each WD
Given fixed user association $\bm{x}$ and RB allocation $\bm{z}$, the original problem (8) can be decomposed into $N$ independent subproblems, each corresponding to the scheduling decision for a single WD. Specifically, for any WD $n\in\mathcal{N}_{a}$ associated with SBS $m$, its local scheduling subproblem can be formulated as:
@@
We define a pair $(c_{n},d_{n})$ as achievable if there exists $f_{n}$ and $P_{n}$ such that the tuple $(c_{n},d_{n},f_{n},P_{n})$ satisfies constraints (9a)-(9c). The set of all achievable workload pairs constitutes the achievable region. Based on our discussion in Section III-B, we can assume the utility function $u^{a}(c_{n},d_{n})$ is non-decreasing, hence the optimal pair $(c^{*}_{n},d^{*}_{n})$ must lie on the boundary of the achievable region, i.e. there is no achievable pair $(c_{n},d_{n})$ such that $c_{n}=c^{*}_{n}$ and $d_{n}>d^{*}_{n}$. To characterize this boundary, it is sufficient to compute the maximum achievable $d_{n}$ for any given $c_{n}$, denoted by $d^{max}_{n}(c_{n})$. This process leads to the following optimization problem:
@@
To solve problem (10), we first demonstrate that its optimal solution satisfies the following tightness condition:
@@
Proof:
We prove by contradiction. We first show that (12) must be satisfied. If $E^{c}_{n}+E^{t}_{n}<E^{max}_{n}$, then we can assign the extra energy to the data transmission process, i.e., let
$P_{n}^{\prime}=P_{n}^{*}+\frac{E^{max}_{n}-E^{c}_{n}-E^{t}_{n}}{T^{t}_{n}}.$
Apparently, $(f^{*}_{n},P_{n}^{\prime})$ is also a feasible solution of problem (10) but produces larger $d_{n}$ than $(f^{*}_{n},P^{*}_{n})$, contradicting the fact that $(f^{*}_{n},P^{*}_{n})$ is optimal.
Next, we show that (11) must be satisfied. If $T^{c}_{n}+T^{t}_{n}<T^{max}$, then we can assign the extra time to the computation process, i.e., let the new computation time be $T^{c,\prime}_{n}=T^{max}-T^{t}_{n}$. To process $c_{n}$ workload in $T^{c,\prime}_{n},$ the corresponding CPU frequency is
$f_{n}^{\prime}=f_{n}^{*}-\frac{(T^{max}-T^{t}_{n}-T^{c}_{n})c_{n}}{T^{c}_{n}(T^{max}-T^{t}_{n})}.$
Since $f_{n}^{\prime}<f_{n}^{*}$, we can save an amount of $\Delta E^{c}_{n}=c_{n}\gamma_{n}[(f_{n}^{*})^{2}-(f_{n}^{\prime})^{2}]$ energy for the computation process. Then the saved energy can be assigned to the communication process to achieve a larger $d_{n}$, i.e. $P_{n}^{\prime}=P^{*}_{n}+\Delta E^{c}_{n}/T^{t}_{n}$. One can easily verify that $(f_{n}^{\prime},P_{n}^{\prime})$ is feasible and produces larger $d_{n}$ than $(f_{n}^{*},P_{n}^{*})$, which results in a contradiction. ∎
With Theorem 2, we can replace the inequality constraints (9a) and (9b) in problem (10) with equalities (11) and (12). Substituting (2) and (6) into (11) and using the fact that WD $n$ is associated with SBS $m$ (i.e. $x^{m}_{n}=1$) result in
@@
Combining (13) with (9c), we can derive the feasible region of $f_{n}$:
@@
Thus, solving problem (10) reduces to finding the optimal $f_{n}$ (denoted as $f_{n}^{*}$) that maximizes (15) within the feasible region defined in (14). This can be achieved by letting $\frac{\partial d_{n}}{\partial f_{n}}=0$ and comparing the values of $d_{n}$ at every critical point in the feasible region. Although the analytical form of the stationary point (where $\frac{\partial d_{n}}{\partial f_{n}}=0$) is intractable due to the complexity of the expression, the value of $f^{*}_{n}$ can be efficiently computed using numerical techniques. [^2] In our implementation, we employ the fzero function in MATLAB, which uses Brent’s method [6], a hybrid root-finding approach combining bisection, secant and inverse quadratic interpolation methods.
Denoting the optimal CPU frequency for a given $c_{n}$ as $f^{*}_{n}(c_{n})$, then we can express $d^{max}_{n}(c_{n})=\max_{f_{n}}d_{n}(c_{n},f_{n})=d_{n}(c_{n},f^{*}_{n}(c_{n}))$, where we regard $d_{n}$ as a function of $c_{n}$ and $f_{n}$ according to (15). Recall that the optimal pair $(c^{*}_{n},d^{*}_{n})$ must lie on the boundary of the achievable region, hence the problem (9) is equivalent to $\max_{c_{n}}u^{a}(c_{n},d^{max}_{n}(c_{n}))$. Similar to $f^{*}_{n}$, the optimal $c^{*}_{n}$ can be obtained by numerically solving
@@
$\quad=\frac{\partial u^{a}}{\partial c_{n}}+\frac{\partial u^{a}}{\partial d^{max}_{n}}\times\frac{\partial d_{n}}{\partial c_{n}}=0,$
where the second equality holds because $d^{max}_{n}(c_{n})=d_{n}(c_{n},f^{*}_{n}(c_{n}))$ and the third equality holds because $\frac{\partial d_{n}}{\partial f^{*}_{n}}=0$. Once $c^{*}_{n}$ is found, the remaining optimal scheduling decisions $f^{*}_{n},P^{*}_{n},d^{*}_{n}$ can be readily computed.
IV-B RB Allocation Subproblem for Each SBS
In the previous subsection, we derived the optimal scheduling decisions $\bm{f},\bm{P},\bm{c},\bm{d}$ under given user association $\bm{x}$ and RB allocation $\bm{z}$. In this subsection, we continue to assume that $\bm{x}$ is fixed and focus on determining the optimal RB allocation $\bm{z}$. Let $u^{*}_{n}(z^{m}_{n})$ denote the maximum utility achieved by WD $n$ (i.e., the optimal objective value of problem (9)) when it is allocated $z^{m}_{n}$ RBs. Given the user association $\bm{x}$, we define $\mathcal{N}_{m}=\{n\in\mathcal{N}\ |\ x^{m}_{n}=1\}$ as the set of WDs associated with SBS $m$. Then the RB allocation subproblem for SBS $m$ can be formulated as:
@@ IV-B RB Allocation Subproblem for Each SBS
In this subsection, we will present two algorithms for problem (16) tailored to different classes of utility function $u^{a}(c,d)$. In many practical applications, such as image recognition and object detection, the utility usually corresponds to inference accuracy, which typically exhibits concavity and other exploitable properties. However, in some scenarios, we may encounter more general utility functions, for which alternative strategies are required.
@@ IV-B1 RB Allocation with Concave Utility
If the utility function represents the inference accuracy or a concave function of the inference accuracy, then it usually satisfies the following three properties. Firstly, according to the results in [43, 41, 42, 7, 31, 46, 47, 20], the marginal returns of both computational and communication workloads are diminishing. Hence, we have:
@@
Secondly, the marginal return of computation workload should be non-decreasing with respect to the amount of transmitted data. This is because any improvement in inference accuracy resulting from additional computation will be diminished if insufficient information is conveyed to the destination. An extreme example is the marginal return $\frac{\partial u^{a}}{\partial c}$ remains zero when $d=0$, as $u^{a}(c,0)=0$ for all $c$. Therefore, we have:
@@
This assumption requires $\frac{\partial u^{a}}{\partial d}$ decreases at least at a linear rate of $d$. A common function that satisfies this property is the logarithmic function. Although Assumption 3 may not hold strictly in all scenarios, it can be regarded as approximately valid within the range of values we are interested in.
Under these assumptions, we can proceed to demonstrate the concavity of $u^{*}_{n}(z^{m}_{n})$. Since the concavity for integer-valued variables is not well-defined, we relax $z^{m}_{n}$ to a continuous variable $\hat{z}^{m}_{n}\in[0,K_{m}]$ and prove that $u^{*}_{n}(\hat{z}^{m}_{n})$ is concave.
@@
Proof:
[…]
$).d^{*}(\hat{z})=\hat{z}g(c^{*}(\hat{z}$)). (18)
[…]
Due to the concavity of $u^{*}_{n}(\hat{z}^{m}_{n})$, we can solve (16) by greedily allocating RBs to the WD that contributes most to the overall utility. The procedure is as follows. Initially, we set $z^{m}_{n}=0$ for all $n\in\mathcal{N}_{m}$. At each iteration, we compute the marginal utility gain for allocating an additional RB to each WD, i.e. $\Delta u_{n}=u^{*}_{n}(z^{m}_{n}+1)-u^{*}_{n}(z^{m}_{n})$. We then identify the WD $n^{\prime}=\operatorname*{arg\,max}_{n\in\mathcal{N}_{m}}\Delta u_{n}$ that yields the highest utility improvement and allocate one extra RB to it, i.e., $z^{m}_{n^{\prime}}=z^{m}_{n^{\prime}}+1$. This process is repeated until all $K_{m}$ RBs are allocated.
Our algorithm is summarized in Algorithm 1. Note that in each iteration, only one variable $z^{m}_{n}$ is updated. Therefore, except for the initial iteration, where all $\Delta u_{n}$ values must be computed, subsequent iterations only require updating $\Delta u_{n}$ for the previously selected WD. The overall time complexity of the algorithm is $O(|\mathcal{N}_{m}|K_{m})$, where $|\mathcal{N}_{m}|$ is the number of elements in $\mathcal{N}_{m}$.
@@ IV-B2 RB Allocation with General Utility
For general utility functions, we cannot leverage the concavity of $u^{*}_{n}(\hat{z}^{m}_{n})$ to facilitate algorithm design. However, the RB allocation subproblem (16) can still be effectively solved using dynamic programming. Specifically, let $U_{m}(k,j)$ be the maximum aggregate utility attainable at SBS $m$ when allocating $k$ RBs among the first $j$ WDs, where $j\in\{1,\dots,|\mathcal{N}_{m}|\}$. Formally, $U_{m}(k,j)$ corresponds to the optimal objective value of the following optimization problem:
@@
Apparently, solving the RB allocation subproblem (16) is equivalent to computing $U_{m}(K_{m},|\mathcal{N}_{m}|)$. For notational convenience, we use $z^{*}_{m}(k,j)$ to denote the number of RBs allocated to the $j$-th WD under the optimal solution to problem (21). Clearly, when there is only one WD in the system, we should allocate all RBs to it. Therefore, we have
$U_{m}(k,1)=u^{*}_{1}(k)\mbox{ and }z^{*}_{m}(k,1)=k.$
For the case with two WDs, we consider all feasible RB allocations $z^{m}_{2}$ to the second WD and assign the remaining $k-z^{m}_{2}$ RBs to the first WD. The optimal allocation is thus given by:
$U_{m}(k,2)=\max_{z^{m}_{2}}\left\{U_{m}(k-z^{m}_{2},1)+u^{*}_{j}(z^{m}_{2})\right\}$
$z^{*}_{m}(k,2)=\operatorname*{arg\,max}_{z^{m}_{2}}\left\{U_{m}(k-z^{m}_{2},1)+u^{*}_{j}(z^{m}_{2})\right\}.$
This recursive formulation can be generalized to any $j\in\{2,3,\dots,N\}$. Specifically, the dynamic programming recursion is given by:
$U_{m}(k,j)=\max_{z^{m}_{j}}\left\{U_{m}(k-z^{m}_{j},j-1)+u^{*}_{j}(z^{m}_{j})\right\}$ (22)
$z^{*}_{m}(k,j)=\operatorname*{arg\,max}_{z^{m}_{j}}\left\{U_{m}(k-z^{m}_{j},j-1)+u^{*}_{j}(z^{m}_{j})\right\}.$ (23)
Based on (22), we can finally obtain $U_{m}(K_{m},|\mathcal{N}_{m}|)$ in a bottom-up manner. After that, the corresponding optimal RB allocation $z^{m}_{n}$ can be reconstructed via backtracking from the recorded $z^{*}_{m}(k,j)$. Specifically, the backtracking procedure begins with $z^{m}_{|\mathcal{N}_{m}|}=z^{*}_{m}(K_{m},|\mathcal{N}_{m}|)$ and for each $j\in\{1,\dots,|\mathcal{N}_{m}|-1\}$, we have
$z^{m}_{j}=z^{*}_{m}\left(K_{m}-\sum_{n=j+1}^{|\mathcal{N}_{m}|}z^{m}_{n},j\right).$
The detailed steps are summarized in Algorithm 2. In each dynamic programming step (22), we iterate over all possible values of $z^{m}_{j}$, which amount to a total of $k$ values. Therefore, the overall computational complexity of Algorithm 2 is $O\left(|\mathcal{N}_{m}|K^{2}_{m}\right)$.
@@
1: Initialization: $U_{m}(k,1)=u^{*}_{1}(k),\ z^{*}_{m}(k,1)=k,\ \forall k\in\{1,\dots,K_{m}\}$;
@@
4:   Calculate $U_{m}(k,j)$ and $z^{*}_{m}(k,j)$ according to (22) and (23).